A spherical segment has a volume of 904.78 . if the radius of the sphere is 10 , solve the area of the zone.
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

the volume formula requires the radius of the top of the segment and the radius of the bottom of the segment and the height of the segment.
you don't have any of them.

the area formula requires the height of the segment.
you don't have that.

the respective formulas are as shown below:

area = 2 * pi * radius of the sphere * height of the segment.

volume = 1/6 * pi * height of the segment * (3 * radius squared of the top of the segment + 3 * radius squared of the bottom of the segment + the height squared of the segment.

the formulas look like this:

A = 2 * pi * R * h

V = 1/6 * pi * h * (3 * R1^2 + 3 * R2^2 + h^2).

hee's a reference.
https://www.geeksforgeeks.org/spherical-segment-formula/

it doesn't appear that i can help you with the information you provided.

i looked at online calculators and none of them seemed to help, given the information you provided.
this thie appears to be due to not enough information provided.

i believe area of the segment referrs to the surface area of the segment which is also called surface area of the zone.

here's some additional information that i found on the web.

https://en.wikipedia.org/wiki/Spherical_segment

you can do a search on "spherical segment" or "spherical segment calculator" to bring up any addiional information from the web.

here's some calculators that i found.

https://rechneronline.de/pi/spherical-segment.php

https://www.redcrab-software.com/en/Calculator/Spherical-Segment

https://www.123calculus.com/en/spherical-segment-page-7-40-140.html

sorry i can't do more.
hopefully this helps.



Answer by ikleyn(52800)

.

The problem is posed incorrectly. As posed, it can not be solved.